if two trials are done and the concentration for trial 1 is 0.004 M and the initial rate for trial 1 is 4.025e-3 and trial 2 is 0.002 M and initial rate is 1.224e-3 how do i calculate the k value for rate law and the n value.
Answer #1

Rate - k [reaclanl] Yo ; rate convlon), n; orden Jrcial 1 ; 4.025 XC-3 -6 10-0041 Trial-2 ; 1: 224 x 3k [o-002] Divide trial( 6-3 0.0498) → 4.025 xe-3-k 0.000016 * k = 251562.580-3 = 251562.5 8000498 = 12527.8 Mesec k= 12527.8 Mivec CAnsieme n=2